OpenStudy (anonymous):

I really need help, what would the answer to this equation be? (sin^2x)(sec^2x) = (sec^2x-1) is this true or not?

OpenStudy (anonymous):

Sec^2 x is 1 over cos^2 x. What does the Tan^2 x =

OpenStudy (anonymous):

Would it be sin x/ cos x^2?

OpenStudy (anonymous):

\[\sec ^2x = \frac{ 1 }{ \cos^2 x }\]

OpenStudy (anonymous):

So yes sin x / cos x = tan x, sin^2 x /cos^2 x = tan^2 x.

OpenStudy (anonymous):

Ok thank you very much, I've been cracking my head for a long time. :)

OpenStudy (anonymous):

Your Welcome! I feel your headache..